Please help me out in solving the following question:
The domain of f(x)= 1/square root(|x|-x) is (-infinite, a) then f(a-18) =________
The options are:
a) 1/6.            b) 1/square root(34).          c) 1/4( square root(2)).            d) 1/8

Grade:11

1 Answers

Aakanksha Talreja
8 years ago
Ans- a) 1/6
\left | x \right |> x \therefore x\epsilon (-infinite,0)
Bcoz it not hold good for x +ve
\therefore f(0-18)=f(-18)=1/\sqrt{36}=1/6